# Quantum Thermoelectric Circuits: A Universal Approach

### v2

@@ Abstract
In this work, we develop a panoramic schematic of a quantum thermoelectric circuit theory in the steady state regime. We establish the foundations of the said premise by defining the analogs of Kirchhoff’s laws for heat currents and temperature gradients. We further show that our approach encompasses various circuits like thermal diode, transistor, and Wheatstone bridge. Additionally, we have been able to develop a model of a quantum thermal step transformer. We also construct a novel model of a thermal adder circuit, paving the way to develop thermal integrated circuits. This sheds new light on the present architecture of quantum device engineering.
@@
Introduction.— All quantum systems are inherently subject to interactions with their surrounding environment, which the theory of open quantum systems addresses [1, 2, 3]. Recent technological advancements have elevated the importance of these interactions, particularly in quantum thermodynamics, where understanding non-equilibrium processes and energy flow at the quantum level has become essential for developing quantum technologies [4, 5, 6, 7, 8, 9].
@@
Logical computational circuits are crucial for the development of quantum computers [10, 11, 12], making quantum circuitry foundational in quantum device engineering. Drawing from classical circuit design, identifying quantum equivalents of diodes, transistors, resistors, and inductors is vital for creating a quantum microprocessor. A promising approach involves using heat currents to create quantum versions of electrical and thermal devices [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Analyzing heat flow patterns in quantum networks is essential to model these quantum thermoelectric devices, focusing on identifying heat current behavior [44, 45]. From the advent of the formal theory of open quantum systems, an important achievement of which was the establishment of the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation [46, 47], modeling quantum heat engines and thermal devices has been one of its most promising applications [48, 49], where heat currents are manipulated to efficient effect. This paper offers a novel quantum thermoelectric network theory, providing the fundamental laws on heat currents and temperature gradients, playing the equivalent roles of electric current and voltage differences, respectively.
@@
We consider the thermoelectric quantum networks from the backdrop of Markovian dynamics [2, 46, 3]. The model involves multiple interacting qubits influenced by non-interacting thermal environmental modes, which can be represented as weak atom-photon couplings, with the baths considered as collections of harmonic oscillators [2]. In this backdrop, the resulting reduced dynamics of the system demonstrates a one-way information flow from the system to the environment, monotonically leading the qubit towards its corresponding thermal equilibrium. It is also important to mention that all our investigations are done in the steady state region, which naturally occurs for such Markovian dynamics. This, in turn, ensures the stability of the circuit models.
@@
We model a quantum thermal resistor identified as the strength of interaction Hamiltonian to find the heat current between two or more qubits. We establish the laws of a thermoelectric circuit, analogous to Kirchhoff’s current and voltage laws, treating a qubit as the node. We use this framework to establish the balance condition of a quantum thermal Wheatstone bridge. In this discourse, we also prove the existence of a quantum thermal step transformer. We further develop a thermal adder circuit, motivated by the operational amplifier circuit, where the voltage output is the algebraic sum of all the inputs. It is also shown that the thermal diode and transistor circuits fall into the framework constructed here, the details of which are presented in the Supplemental Material [50]. The framework introduced here is the first of its kind, providing a fundamental tool to build a comprehensive quantum thermoelectric circuit theory.
@@
Quantum thermal resistor and current law.— Our program begins by constructing a model for a quantum thermal resistor. By analyzing the flow of heat current through a quantum system between two thermal baths with differing temperatures, we seek to establish a relationship akin to Kirchhoff’s laws. Our goal is to understand if, and under what conditions, temperature gradients and heat currents in quantum thermal devices can relate in a manner analogous to these fundamental circuit principles. The quantum thermal resistor is made up of two qubits, Fig. 1(a).
@@
Figure 1: Schematic diagrams of the (a) quantum thermal resistor and (b) 3-qubit quantum thermal circuit. Quantum thermal current from source $j$ to qubit $k$, taking qubit $k$ as a node, is $\mathcal{J}_{jk}$.
@@
Qubits 1 and 2 are under the influence of bosonic baths at temperatures zero and $T$, respectively. The Hamiltonian of the two-qubit system (for $\hbar=k_{B}=1$) is given by $H_{S}=\sum_{j=1}^{2}\frac{\omega_{j}}{2}\sigma_{j}^{z}+J_{12}\left(\sigma^{x}_{1}\sigma^{x}_{2}+\sigma^{y}_{1}\sigma^{y}_{2}\right),$ where $\sigma^{i}~{}(i=x,y,z)$ are the Pauli matrices. $J_{12}$ is the interaction strength between the two qubits and models the quantum thermal resistor. The dynamics of the system is governed by the GKSL master equation of the form
@@
where the forms of $\alpha_{ij}$ and $\eta_{ij}$ are given in [50] and |$\ket{i}$ is in the computational basis. The quantum heat currents, in the steady state condition, from the bath $j$ to the qubit $k$ ($\mathcal{J}_{jk}$) and from qubit $l$ to qubit $k$ ($\mathcal{J}_{lk}$) are given by
@@
respectively [50].
Using the state $\rho^{SS}$ from Eq. (2), the current $\mathcal{J}_{A1}$ from bath $A$ to qubit 1 is $\mathcal{J}_{A1}={\rm Tr}\left[H_{1}\mathcal{D}_{A1}\left(\rho^{SS}\right)\right]=-\gamma_{A}\omega_{1}\left(\frac{\alpha_{00}}{\eta_{00}}+\frac{\alpha_{11}}{\eta_{11}}\right).$ Pertinently, since the parameters $\gamma_{A}\geq 0$, $\alpha_{00}/\eta_{00}\geq 0$ and $\alpha_{11}/\eta_{11}\geq 0$, the current $\mathcal{J}_{A1}$ negative. The rationale behind $\mathcal{J}_{A1}$ being negative is that qubit 1 is surrounded by a zero temperature bath, and negative $\mathcal{J}_{A1}$ indicates that the heat flow is directed from qubit 1 towards the thermal reservoir. Therefore, reservoir $A$ acts as a “ground” to which the quantum thermal resistor is connected. Furthermore, the current $\mathcal{J}_{21}$ from qubit 2 to 1 is given by $\mathcal{J}_{21}=\frac{2iJ_{12}\omega_{1}\left(\alpha_{12}-\alpha_{12}^{*}\right)}{\eta_{12}}=\frac{-4J_{12}\omega_{1}\Im\left(\alpha_{12}\right)}{\eta_{12}},$ where $\Im(*)$ denotes imaginary part. Interestingly, we find $\mathcal{J}_{A1}=-\left(\frac{\gamma_{A}\omega_{1}x_{0}}{\gamma_{B}}\right)\frac{\alpha_{00}}{\eta_{00}}=-\mathcal{J}_{21}$, resulting in $\mathcal{J}_{A1}+\mathcal{J}_{21}=0.$ This is crucial as it allows qubit 1 to be considered as a junction node, leading to the visualization of the above equation as the quantum thermal version of Kirchhoff’s current law. In fact, considering heat currents $\mathcal{J}_{jk}$’s from $n$ different sources to a qubit $k$, it can be proved that in the steady state condition (see [50]), the sum of all the heat currents is zero, i.e.,
@@
In a similar manner, we find that the currents with qubit 2 as the junction node are given by $\mathcal{J}_{B2}={\rm Tr}\left[H_{2}\mathcal{D}_{B2}\left(\rho^{SS}\right)\right]=\left(\frac{\gamma_{A}\omega_{2}x_{0}}{\gamma_{B}}\right)\frac{\alpha_{00}}{\eta_{00}}$, and $\mathcal{J}_{12}=-\left(\frac{\gamma_{A}\omega_{2}x_{0}}{\gamma_{B}}\right)\frac{\alpha_{00}}{\eta_{00}}.$ It can be easily verified that $\mathcal{J}_{B2}+\mathcal{J}_{12}=0$, which, upon assuming qubit 2 as a junction node, shows the characteristics of the quantum thermal version of Kirchoff’s current law.
Quantum thermal transformer.— Interestingly, we observe that the heat currents $\mathcal{J}_{21}$ and $\mathcal{J}_{12}$ are interconnected by the relation $\frac{\mathcal{J}_{21}}{\omega_{1}}=-\frac{\mathcal{J}_{12}}{\omega_{2}},$ or $\left|\frac{\mathcal{J}_{21}}{\mathcal{J}_{12}}\right|=\frac{\omega_{1}}{\omega_{2}}=\left|\frac{\mathcal{J}_{A1}}{\mathcal{J}_{B2}}\right|.$ In general, for an interaction $H_{jk}=J_{jk}\left(\sigma^{x}_{j}\sigma^{x}_{k}+\sigma^{y}_{j}\sigma^{y}_{k}\right)$ between two qubits $j$ and $k$, the current between them can be shown to be related by
@@
The proof is given in [50]. The above relation mimics the quantum thermal version of a step transformer from electrical circuits, where the ratio of electric currents between two coils depends on the number of turns in each coil. Here, the ratio of quantum heat currents depends on the transition frequencies $\omega_{j}$ and $\omega_{k}$ of the qubits $j$ and $k$, respectively, thereby allowing for the manipulation of quantum heat currents.
Quantum thermal voltage.— We now develop the quantum analog of the thermal voltage driving the thermal current. To this end, we determine the effective temperature of the qubits 1 and 2 in the steady state. The reduced state of qubit 1 from the steady state, Eq. 2, is given by $\rho^{SS}_{1}=\left(\rho^{SS}_{00}+\rho^{SS}_{11}\right)\ket{0}\bra{0}+\left(\rho^{SS}_{22}+\rho^{SS}_{33}\right)\ket{1}\bra{1}.$ | . The effective temperature $T_{1}$ of the qubit 1 is found by comparing the state $\rho^{SS}_{1}$ with the state $e^{-H_{1}/T_{1}}/{\rm Tr}\left(e^{-H_{1}/T_{1}}\right)$, where $H_{1}=\frac{\omega_{1}}{2}\sigma^{z}_{1}$, leading to $T_{1}=\omega_{1}/\log\left(\frac{1}{\rho^{SS}_{00}+\rho^{SS}_{11}}-1\right)=\omega_{1}/\log\left(\frac{\gamma_{B}\eta_{00}}{x_{0}\alpha_{00}}-1\right).$ Note that $T_{1}$ can be envisaged as the thermal potential difference between the qubit 1 and bath $A$, that is, $V_{1A}=T_{1}-0=T_{1}$. This potential difference drives the thermal current $\mathcal{J}_{A1}$, which explicitly in terms of $T_{1}$ can be written as $\mathcal{J}_{A1}=\frac{-\gamma_{A}\omega_{1}}{1+e^{\omega_{1}/T_{1}}}.$ In the low and high-temperature limits, this is
$\mathcal{J}_{A1}=\begin{cases}\frac{-\gamma_{A}\omega_{1}}{2},&\text{for}~{}~{}T_{1}\gg 0\\ -\gamma_{A}\omega_{1}e^{-\omega_{1}/T_{1}}.&\text{for}~{}~{}T_{1}\sim 0\end{cases}$ (6)
@@
This has been depicted in Fig. 2. Similarly, one can find the reduced state of the qubit 2 and its corresponding effective temperature, which is $T_{2}=\omega_{2}/\log\left(\frac{1}{\rho^{SS}_{00}+\rho^{SS}_{22}}-1\right).$ The potential difference between the bath $B$ and the qubit 2 can now be written as $V_{B2}=T-T_{2}$. Further, $\mathcal{J}_{B2}=\frac{\gamma_{B}\omega_{2}}{2}\left[-1+\coth\left(\frac{\omega_{2}}{2T}\right)\tanh\left(\frac{\omega_{2}}{2\left\{T-V_{B2}\right\}}\right)\right].$ It can be pointed out here that the thermal current $\mathcal{J}_{B2}$ does not explicitly depend only on the thermal potential difference $V_{B2}$ but also on the temperature ($T$) of the bath $B$.
@@
The heat currents $\mathcal{J}_{jk}$ discussed here are plotted in Fig. 3 as a function of temperature $T$. The heat current can be seen to increase linearly in certain parameter regimes but, in general, exhibits non-linear behavior. However, Kirchhoff’s current law is satisfied universally.
Law of thermal potentials.— Having discussed quantum thermal resistors, we now move on to study quantum thermal circuits made up of multiple qubits. For this purpose, we consider quantum thermal circuits composed of three qubits. A schematic diagram of this circuit is given in Fig.  1(b), where qubits 1 and 2 are under the influence of bosonic baths $A$ and $B$ at zero and $T$ temperatures, respectively. In this setup, we consider a third qubit that interacts with qubits 1 and 2, which themselves interact with each other. The Hamiltonian for this three-qubit setup is $\widetilde{H}_{S}=\frac{1}{2}\sum_{i=1}^{3}\omega_{i}\sigma^{z}_{i}+\sum_{l,k=1,l<k}^{3}J_{lk}\left(\sigma^{x}_{l}\sigma^{x}_{k}+\sigma^{y}_{l}\sigma^{y}_{k}\right),$ where qubits $l$ and $k$ interact with each other via Heisenberg $XX$ type interaction with the interaction strength $J_{lk}$ and $\omega_{k}$ being the transition frequency of the $k$-th qubit. We compute the steady-state $\tilde{\rho}^{SS}$ of this system using Eq. (1) by replacing $H_{S}$ with $\widetilde{H}_{S}$.
@@
Figure 5: Schematic diagrams of the (a) quantum thermal Wheatstone Bridge and (b) quantum thermal adder circuit.
The variation of the quantum heat currents in this circuit is shown in Fig. 4. It can be verified from this figure that Kirchhoff’s current law, which is discussed above, holds true for quantum thermal currents at a junction node. Further, the effective temperatures $T_{1}$, $T_{2}$, and $T_{3}$ of the qubits 1, 2, and 3, respectively, can be calculated using the steady-state $\tilde{\rho}^{SS}$ of the three-qubit system similarly to the previous case, and are given by $T_{1}=\omega_{1}/\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{11}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{33}}-1\right)$, $T_{2}=\omega_{2}/\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{11}+\tilde{\rho}^{SS}_{44}+\tilde{\rho}^{SS}_{55}}-1\right)$, and $T_{3}=\omega_{3}/\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{44}+\tilde{\rho}^{SS}_{66}}-1\right)$. The corresponding thermal potential differences are given by $V_{1A}=T_{1}-0;V_{21}=T_{2}-T_{1};V_{32}=T_{3}-T_{2};V_{13}=T_{1}-T_{3};V_{B2}=T-T_{2}.$ Using this, we get a relation for the potential differences inside the loop of the three qubits, which is $V_{21}+V_{32}+V_{13}=0$, resembling Kirchhoff’s second law of an electric circuit, which states that the sum of potential differences around a closed loop in an electric circuit is zero. The relation between the thermal current $\mathcal{J}_{A1},\mathcal{J}_{B2}$ and the potential difference $V_{1A},V_{B2}$ is the same as in the case of the previous circuit. In terms of the elements of the steady state $\tilde{\rho}^{SS}$ of the three-qubit circuit, the quantum thermal currents from qubit 3 to 1 and from qubit 2 to 1 are given by $\mathcal{J}_{31}=-4J_{13}\omega_{1}\left[\Im{(\tilde{\rho}^{SS}_{14})}+\Im{(\tilde{\rho}^{SS}_{36})}\right]$ and $\mathcal{J}_{21}=-4J_{12}\omega_{1}\left[\Im{(\tilde{\rho}^{SS}_{24})}+\Im{(\tilde{\rho}^{SS}_{35})}\right]$, respectively. This leads to a relation between the quantum thermal heat current $\mathcal{J}_{31}$ ($\mathcal{J}_{13}$) and the thermal potential $V_{13}$ given by

### v3

# Quantum Thermal Analogs of Electric Circuits: A Universal Approach

@@ Abstract
In this work, we develop a panoramic schematic of quantum thermal analogs of electric circuits in the steady state regime. We establish the foundations of said premise by defining the analogs of Kirchhoff’s laws for heat currents and temperature gradients, as well as a quantum thermal step transformer. Using this, we develop two novel quantum thermal circuits, viz., quantum thermal super Wheatstone bridge and quantum thermal adder circuit, paving the way for the corresponding integrated circuits. We further show that our approach encompasses various circuits like thermal diode, transistor, and Wheatstone bridge. This sheds new light on the present architecture of quantum device engineering.
@@
Introduction—All quantum systems are inherently subject to interactions with their surrounding environment, which the theory of open quantum systems addresses [1, 2, 3]. Recent technological advancements have elevated the importance of these interactions, particularly in quantum thermodynamics, where understanding non-equilibrium processes and energy flow at the quantum level has become essential for developing quantum technologies [4, 5, 6, 7, 8, 9].
@@
Logical computational circuits are crucial for the development of quantum computers [10, 11, 12], making quantum circuitry foundational in quantum device engineering. Drawing from classical circuit design, identifying quantum equivalents of diodes, transistors, resistors, and inductors is vital for creating a quantum microprocessor. A promising approach involves using heat currents to create quantum versions of electrical and thermal devices [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Analyzing heat flow patterns in quantum networks is essential to model these quantum thermoelectric devices, focusing on identifying heat current behavior [48, 49]. From the advent of the formal theory of open quantum systems, an important achievement of which was the establishment of the Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) master equation [50, 51], modeling quantum thermal devices has been one of its most promising applications [52, 53], where heat currents are manipulated to efficient effect. This paper offers a novel quantum thermoelectric network theory, providing the fundamental laws on heat currents and temperature gradients, playing the equivalent roles of electric current and voltage differences, respectively.
@@
We consider the thermoelectric quantum networks from the backdrop of Markovian dynamics [2, 50, 3]. The model involves multiple interacting qubits influenced by non-interacting thermal environmental modes, with the baths considered as collections of harmonic oscillators [2]. In this backdrop, the resulting reduced dynamics of the system demonstrates a one-way information flow from the system to the environment, monotonically leading the qubit toward its corresponding thermal equilibrium. It is also important to mention that all our investigations are done in the steady state region, which naturally occurs for such Markovian dynamics. This, in turn, ensures the stability of the circuit models.
@@
We model a quantum thermal resistor identified as the strength of interaction Hamiltonian to find the heat current between two or more qubits. We establish the laws of a quantum thermal analog of an electric circuit, analogous to Kirchhoff’s current and voltage laws, treating a qubit as the node. We also prove the existence of a quantum thermal step transformer. We develop two novel quantum thermal circuits—quantum thermal super Wheatstone bridge, and a thermal adder circuit (motivated by the operational amplifier circuit, where the voltage output is the algebraic sum of all the inputs)—and use the above framework to study them. It is shown that the thermal diode, transistor, and the standard Wheatstone bridge circuits fall into the framework constructed here, the details of which are presented in Supplemental Material [54]. The framework introduced is the first of its kind, providing a fundamental tool to build a comprehensive quantum thermal circuit theory. The fact that many thermal versions of important classical electric circuits can be studied in a unified manner under the umbrella of this framework will interest a broader audience in the juxtaposition of quantum thermodynamics and electric circuits.
@@
Quantum thermal resistor and current law—Our program begins by constructing a model for a quantum thermal resistor. By analyzing the flow of heat current through a quantum system between two thermal baths with differing temperatures, we seek to establish a relationship akin to Kirchhoff’s laws. Our goal is to understand if, and under what conditions, temperature gradients and heat currents in quantum thermal devices can relate in a manner analogous to these fundamental circuit principles. The quantum thermal resistor is made up of two qubits, Fig. 1(a).
@@
Figure 1: Schematic diagrams of the (a) quantum thermal resistor and (b) three-qubit quantum thermal circuit. The quantum thermal current from source $j$ to qubit $k$, taking qubit $k$ as a node, is $\mathcal{J}_{jk}$.
@@
Qubits 1 and 2 are under the influence of bosonic baths at temperatures zero and $T$, respectively. The Hamiltonian of this system (for $\hbar=k_{B}=1$) is $H_{S}=\sum_{j=1}^{2}\frac{\omega_{j}}{2}\sigma_{j}^{z}+J_{12}\left(\sigma^{x}_{1}\sigma^{x}_{2}+\sigma^{y}_{1}\sigma^{y}_{2}\right),$ where $\sigma^{i}~{}(i=x,y,z)$ are the Pauli matrices. $J_{12}$ is the interaction strength between the two qubits and models the quantum thermal resistor. The dynamics of the system is governed by
@@
[…]
@@
respectively [54]. The above heat currents are derived from the time derivative of heat transferred to the qubit, providing heat current between individual components of the circuit. Considering a spin-chain, the above heat current formula becomes equivalent to the total heat current ${\rm Tr}\left[H_{S}\mathcal{D}_{jk}(\rho^{SS})\right]$ for equal qubit frequencies $\omega_{j}$ (see [54]), which is fundamental to studies in thermal conductivity [53, 55]. Further, the heat current between the nodes is also related to the spin current [56, 57, 58, 59], as shown in [54].
Using the state $\rho^{SS}$ from Eq. (2), the current $\mathcal{J}_{A1}$ from bath $A$ to qubit 1 is $\mathcal{J}_{A1}={\rm Tr}\left[H_{1}\mathcal{D}_{A1}\left(\rho^{SS}\right)\right]=-\gamma_{A}\omega_{1}\left(\frac{\alpha_{00}}{\eta_{00}}+\frac{\alpha_{11}}{\eta_{11}}\right).$ Pertinently, since the parameters $\gamma_{A}\geq 0$, $\alpha_{00}/\eta_{00}\geq 0$ and $\alpha_{11}/\eta_{11}\geq 0$, the current $\mathcal{J}_{A1}$ negative. The rationale behind $\mathcal{J}_{A1}$ being negative is that qubit 1 is surrounded by a zero-temperature bath, and negative $\mathcal{J}_{A1}$ indicates that the heat flow is directed from qubit 1 toward the thermal reservoir. Therefore, reservoir $A$ acts as a “ground” to which the quantum thermal resistor is connected. Furthermore, the current $\mathcal{J}_{21}$ from qubit 2 to 1 is given by $\mathcal{J}_{21}=\frac{2iJ_{12}\omega_{1}\left(\alpha_{12}-\alpha_{12}^{*}\right)}{\eta_{12}}=\frac{-4J_{12}\omega_{1}\Im\left(\alpha_{12}\right)}{\eta_{12}},$ where $\Im(*)$ denotes imaginary part. Interestingly, we find $\mathcal{J}_{A1}=-\left(\frac{\gamma_{A}\omega_{1}x_{0}}{\gamma_{B}}\right)\frac{\alpha_{00}}{\eta_{00}}=-\mathcal{J}_{21}$, resulting in $\mathcal{J}_{A1}+\mathcal{J}_{21}=0.$ This is crucial as it allows qubit 1 to be considered as a junction node, leading to the visualization of the above equation as the quantum thermal version of Kirchhoff’s current law. In fact, considering heat currents $\mathcal{J}_{jk}$’s from $n$ different sources to a qubit $k$, it can be proved that in the steady state condition (see [54]), the sum of all the heat currents is zero, i.e.,
@@
Similarly, we find that the currents with qubit 2 as the junction node are given by $\mathcal{J}_{B2}={\rm Tr}\left[H_{2}\mathcal{D}_{B2}\left(\rho^{SS}\right)\right]=\left(\frac{\gamma_{A}\omega_{2}x_{0}}{\gamma_{B}}\right)\frac{\alpha_{00}}{\eta_{00}}$, and $\mathcal{J}_{12}=-\left(\frac{\gamma_{A}\omega_{2}x_{0}}{\gamma_{B}}\right)\frac{\alpha_{00}}{\eta_{00}}.$ It can be easily verified that $\mathcal{J}_{B2}+\mathcal{J}_{12}=0$, which shows the characteristics of the quantum thermal version of Kirchhoff’s current law.
Quantum thermal transformer.— Interestingly, the heat currents $\mathcal{J}_{21}$ and $\mathcal{J}_{12}$ are interconnected by the relation $\frac{\mathcal{J}_{21}}{\omega_{1}}=-\frac{\mathcal{J}_{12}}{\omega_{2}},$ or $\left|\frac{\mathcal{J}_{21}}{\mathcal{J}_{12}}\right|=\frac{\omega_{1}}{\omega_{2}}=\left|\frac{\mathcal{J}_{A1}}{\mathcal{J}_{B2}}\right|.$ In general, for an interaction $H_{jk}=J_{jk}\left(\sigma^{x}_{j}\sigma^{x}_{k}+\sigma^{y}_{j}\sigma^{y}_{k}\right)$ between two qubits $j$ and $k$, the current between them can be shown to be [54]
@@
The above relation mimics the quantum thermal version of a step transformer, where the ratio of currents between two coils depends on the number of turns in each coil. Here, the ratio of quantum heat currents depends on the transition frequencies $\omega_{j}$ and $\omega_{k}$ of the qubits $j$ and $k$, respectively, thereby allowing their manipulation.
Quantum thermal voltage—We now develop the quantum analog of the thermal voltage driving the thermal current. To this end, we determine the effective temperature of qubits 1 and 2 in the steady state. The reduced state of qubit 1 from the steady state, Eq. 2, is given by $\rho^{SS}_{1}=\left(\rho^{SS}_{00}+\rho^{SS}_{11}\right)\ket{0}\bra{0}+\left(\rho^{SS}_{22}+\rho^{SS}_{33}\right)\ket{1}\bra{1}.$ | . The effective temperature $T_{1}$ of the qubit 1 is found by comparing the state $\rho^{SS}_{1}$ with the state $e^{-H_{1}/T_{1}}/{\rm Tr}\left(e^{-H_{1}/T_{1}}\right)$, where $H_{1}=\frac{\omega_{1}}{2}\sigma^{z}_{1}$, leading to $T_{1}=\omega_{1}/\log\left(\frac{1}{\rho^{SS}_{00}+\rho^{SS}_{11}}-1\right)=\omega_{1}/\log\left(\frac{\gamma_{B}\eta_{00}}{x_{0}\alpha_{00}}-1\right).$ Note that $T_{1}$ can be envisaged as the thermal potential difference between qubit 1 and bath $A$, that is, $V_{1A}=T_{1}-0=T_{1}$. This potential difference drives the thermal current $\mathcal{J}_{A1}$, which explicitly in terms of $T_{1}$ can be written as $\mathcal{J}_{A1}=\frac{-\gamma_{A}\omega_{1}}{1+e^{\omega_{1}/T_{1}}}.$ In the low and high-temperature limits, this is
$\mathcal{J}_{A1}=\begin{cases}\frac{-\gamma_{A}\omega_{1}}{2}&\text{for}~{}~{}T_{1}\gg 0,\\ -\gamma_{A}\omega_{1}e^{-\omega_{1}/T_{1}}&\text{for}~{}~{}T_{1}\sim 0.\end{cases}$ (6)
@@
This has been depicted in Fig. 2. Similarly, one can find the reduced state of qubit 2 and its corresponding effective temperature, which is $T_{2}=\omega_{2}/\log\left(\frac{1}{\rho^{SS}_{00}+\rho^{SS}_{22}}-1\right).$ The potential difference between the bath $B$ and the qubit 2 can now be written as $V_{B2}=T-T_{2}$. Further, $\mathcal{J}_{B2}=\frac{\gamma_{B}\omega_{2}}{2}\left[-1+\coth\left(\frac{\omega_{2}}{2T}\right)\tanh\left(\frac{\omega_{2}}{2\left\{T-V_{B2}\right\}}\right)\right].$ It can be pointed out here that the thermal current $\mathcal{J}_{B2}$ does not explicitly depend only on the thermal potential difference $V_{B2}$ but also on the temperature ($T$) of the bath $B$.
@@
The heat currents $\mathcal{J}_{jk}$ discussed here are plotted in Fig. 3 as a function of temperature $T$. The heat current can be seen to increase linearly in certain parameter regimes. Thus, for example, in Fig. 3(d), for $\omega_{2}\gg\omega_{1}$, we find the current $\mathcal{J}_{21}$ increases almost linearly with $T$ (particularly, for $J_{12}=0.5$ and $J_{12}=1$), satisfying the condition for Ohm’s law. However, in general, the heat current exhibits nonlinear behavior. Further, Kirchhoff’s current law is satisfied universally.
Law of thermal potentials—We now move on to study quantum thermal circuits made up of multiple qubits. A schematic diagram is given in Fig.  1(b), with the Hamiltonian $\widetilde{H}_{S}=\frac{1}{2}\sum_{i=1}^{3}\omega_{i}\sigma^{z}_{i}+\sum_{l,k=1,l<k}^{3}J_{lk}\left(\sigma^{x}_{l}\sigma^{x}_{k}+\sigma^{y}_{l}\sigma^{y}_{k}\right)$, where qubits $l$ and $k$ interact with each other with the interaction strength $J_{lk}$ and $\omega_{k}$ being the transition frequency of the $k$-th qubit. Qubits 1 and 2 are under the influence of the bath. We compute the steady-state $\tilde{\rho}^{SS}$ of this system using Eq. (1) by replacing $H_{S}$ with $\widetilde{H}_{S}$.
@@
The variation of the quantum heat currents in this circuit is shown in Fig. 4. It can be verified from this figure that Kirchhoff’s current law holds true at a junction node. Further, the effective temperatures of the qubits 1, 2, and 3 are given by $T_{1}=\omega_{1}/\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{11}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{33}}-1\right)$, $T_{2}=\omega_{2}/\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{11}+\tilde{\rho}^{SS}_{44}+\tilde{\rho}^{SS}_{55}}-1\right)$, and $T_{3}=\omega_{3}/\log\left(\frac{1}{\tilde{\rho}^{SS}_{00}+\tilde{\rho}^{SS}_{22}+\tilde{\rho}^{SS}_{44}+\tilde{\rho}^{SS}_{66}}-1\right)$, respectively. The corresponding thermal potential differences are $V_{1A}=T_{1}-0;V_{21}=T_{2}-T_{1};V_{32}=T_{3}-T_{2};V_{13}=T_{1}-T_{3};V_{B2}=T-T_{2}.$
Figure 5: A schematic diagram of the quantum thermal super Wheatstone Bridge.
Using this, we get a relation $V_{21}+V_{32}+V_{13}=0$ inside the loop of the three qubits, resembling Kirchhoff’s second law of an electric circuit, stating that the sum of potential differences around a closed loop (a closed sequence of interactions connecting qubits, returning to the initial qubit) in an electric circuit is zero. The relation between the thermal currents $\mathcal{J}_{A1},\mathcal{J}_{B2}$ and the potential differences $V_{1A},V_{B2}$ is the same as in the case of the previous circuit. The currents from qubit 3 to 1 and from qubit 2 to 1 are given by $\mathcal{J}_{31}=-4J_{13}\omega_{1}\left[\Im{(\tilde{\rho}^{SS}_{14})}+\Im{(\tilde{\rho}^{SS}_{36})}\right]$ and $\mathcal{J}_{21}=-4J_{12}\omega_{1}\left[\Im{(\tilde{\rho}^{SS}_{24})}+\Im{(\tilde{\rho}^{SS}_{35})}\right]$, respectively. This leads to the relation [54]